OpenStudy (anonymous):

Four friends have different numbers of cows on their farms. Jorge has 3 times as many cows as Carla. Marcus has half as many cows as Jorge. Dominique has 2 less than 4 times as many cows as Carla. How many cows does each person have if they have a total of 169 cows?

OpenStudy (anonymous):

x: cows of Carla 169 = 3 x + 3/2 x + 4x-2 + x total = jorge + Marcus + Dominique + Carla

OpenStudy (anonymous):

169 = 6/2 x + 3/2 x + 8/2 x -2 + 2/2 x 169 = 9/2 x + 10/2 x -2 169 = 19/2 x -2 171 = 19/2 x

OpenStudy (anonymous):

19 x = 171 x 2 19 x = 342 x = 18 according to the calculations, carla has 18 cows

OpenStudy (anonymous):

thank you so much i am a old lady who has a grandson who is having trouble with algerbra and i have no clue how to help him